Find two numbers such that the first with half the second may make 55, and the second diminished by a fourth of the first may give 20.

Answer by Cromlix(4381)

Hi there,
the first with half the second may make 55.
x + y/2 = 55.......(1)
the second diminished by a fourth of the first may give 20.
y - x/4 = 20.......(2)
...........
x + y/2 = 55.......(1)
y - x/4 = 20.......(2)
Rearrange Eq(1)
y/2 + x = 55.......(1)
y - x/4 = 20.......(2)
Multiply Eq(2) by 4
y/2 + x = 55.......(1)
4y - x = 80.......(2)
Add Eq(1) + Eq(2)
4.5y = 135
y = 30
Substitute y = 30 into Eq(1).
x + y/2 = 55.......(1)
x + 30/2 = 55
x + 15 = 55
x = 55 - 15
x = 40
First number = 40
Second number = 30
Hope this helps :-)